# Factorized Tensor Networks for Multi-Task and Multi-Domain Learning

### v1

@@ Abstract
Multi-task and multi-domain learning methods seek to learn multiple tasks/domains, jointly or one after another, using a single unified network. The key challenge and opportunity is to exploit shared information across tasks and domains to improve the efficiency of the unified network. The efficiency can be in terms of accuracy, storage cost, computation, or sample complexity. In this paper, we propose a factorized tensor network (FTN) that can achieve accuracy comparable to independent single-task/domain networks with a small number of additional parameters. FTN uses a frozen backbone network from a source model and incrementally adds task/domain-specific low-rank tensor factors to the shared frozen network. This approach can adapt to a large number of target domains and tasks without catastrophic forgetting. Furthermore, FTN requires a significantly smaller number of task-specific parameters compared to existing methods. We performed experiments on widely used multi-domain and multi-task datasets. We show the experiments on convolutional-based architecture with different backbones and on transformer-based architecture. We observed that FTN achieves similar accuracy as single-task/domain methods while using only a fraction of additional parameters per task.
@@ 1 Introduction
The primary objective in multi-task learning (MTL) is to train a single model to learn multiple related tasks, either jointly or sequentially. Multi-domain learning (MDL) aims to achieve the same learning objective across multiple domains. MTL and MDL techniques seek to improve overall performance by leveraging shared information across multiple tasks and domains. On the other hand, single-task or single-domain learning do not have that opportunity. Furthermore, the storage and computational cost associated with single-task/domain models quickly grows as the number of tasks/domains increases. In contrast, MTL and MDL methods can use the same network resources for multiple tasks/domains, which keeps the overall computational and storage cost small [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
@@
where $\mathbf{F}_{t}$ represents a function for task/domain $t$ that maps input $\mathbf{x}$ to output $\mathbf{y}_{t}$. We further assume that $\mathbf{F}$ represents a network with a fixed architecture and $\mathcal{W}_{t}$ and $h_{t}$ represent the parameters for task/domain-specific feature extraction and classification/inference heads, respectively. The function in (1) can represent the network for specific task/domain $t$ using the respective $\mathcal{W}_{t},h_{t}$. In the case of MTL, with $T$ tasks, we can have $T$ outputs $\mathbf{y}_{1},\ldots,\mathbf{y}_{T}$ for a given input $\mathbf{x}$. In the case of MDL, we usually have a single output for a given input, conditioned on the domain $t$. Our main goal is to learn the $\mathcal{W}_{t},h_{t}$ for all $t$ that maximize the performance of MTL/MDL with minimal computation and memory overhead compared to single task/domain learning.
@@
Figure 1(a),(b),(c) illustrate three typical approaches for MTL/MDL. First, we can start with a pre-trained network and fine-tune all the parameters $(\mathcal{W}_{t})$ to learn a target task/domain, as shown in Figure 1(a). Fine-Tuning approaches can transfer some knowledge from the pretrained network to the target task/domain, but they effectively use an independent network for every task/domain [1, 5, 11, 12, 13, 14]. Second, we can reduce the parameter and computation complexity by using a completely shared Feature-Extractor (i.e., $\mathcal{W}_{t}=\mathcal{W}_{\text{shared}}$ for all $t$) and learning task/domain-specific heads as last layers, as shown in Figure 1(b). While such approaches reduce the number of parameters, they often result in poor overall performance because of limited network capacity and interference among features for different tasks/domains [1, 4, 5, 15]. Third, we can divide the network into shared and task/domain-specific parameters or pathways, as shown in Figure 1(c). Such an approach can increase the network capacity, provide interference-free paths for task/domain-specific feature extraction, and enable knowledge sharing across the tasks/domains. In recent years, a number of such methods have been proposed for MTL/MDL [1, 4, 5, 9, 16, 17, 18, 19, 20, 21, 22]. While existing methods provide performance comparable to single-task/domain learning, they require a significantly large number of parameters.
In this paper, we propose a new method to divide network into shared and task/domain-specific parameters using a factorized tensor network (FTN). In particular, our method learns task/domain-specific low-rank tensor factors and normalization layers. An illustration of our proposed method is shown in Figure 1(d), where we represent network parameters as $\mathcal{W}_{t}=\mathcal{W}_{\text{shared}}+\Delta\mathcal{W}_{t}$, where $\Delta\mathcal{W}_{t}$ is a low-rank tensor. Furthermore, we also learn task/domain-specific normalization parameters. We demonstrate the effectiveness of our method using different MTL and MDL datasets. Our method can achieve accuracy comparable to a single-task/domain network with a small number of additional parameters. Existing parameter-efficient MTL/MDL methods [1, 2, 23] introduce small task/domain-specific parameters while others [15, 24] add many parameters to boost the performance irrespective of the task complexity. In our work, we demonstrate the flexibility of FTNs by selecting the rank according to the complexity of the task.
@@
The main contributions of this paper can be summarized as follows.
  • We propose a new method for MTL and MDL, called factorized tensor networks (FTN), that adds task/domain-specific low-rank tensors to shared weights.
  • We demonstrate that by using a fraction of additional parameters per task/domain, FTNs can achieve similar performance as the single-task/domain methods.
@@
  • We performed empirical analysis to show that the FTNs enable flexibility by allowing us to vary the rank of the task-specific tensors according to the problem complexity.
Limitations.
In our experiments, we used a fixed rank for each layer. In principle, we can adaptively select the rank for different layers to further reduce the parameters. MTL/MDL models often suffer from task interference or negative transfer learning when multiple conflicting tasks are trained jointly. Our method can have similar drawbacks as we did not investigate which tasks/domains should be learned jointly. A shared backbone requires a single forward pass for all tasks while our proposed FTN would require as many forward passes as the number of tasks. Branched and tree-structures can enable different tasks to share several layers and reduce latency.
@@ 2 Related Work
Multi-task learning (MTL) methods commonly leverage shared and task-specific layers in a unified network to solve related tasks [16, 17, 18, 25, 26, 27, 28, 29, 30, 31, 32, 33]. These methods learn shared and task-specific representation through their respective modules. Optimization based methods [34, 35, 36, 37, 38] devise a principled way to evaluate gradients and losses in multi-task settings. Branched and tree-structured MTL methods [27, 39, 40] enable different tasks to share branches along a tree structure for several layers. Multiple tasks can share computations and features in any layer only if they belong to the same branch in all the preceding layers. MTL networks that incrementally learn new taks were proposed in [9, 10]. ASTMT [10] proposed a network that emphasizes or suppresses features depending on the task at hand. RCM [9] reparameterizes the convolutional layer into non-trainable and task-specific trainable modules. We compare our proposed method with these incrementally learned networks. Adashare [8] is another related work in MTL that jointly learns multiple tasks. It learns task-specific polices and network pathways [41].joint learning setups.
[…]
@@
Domain adaptation and transfer learning. The work in this field usually focuses on learning a network from a given source domain to a closely related target domain. The target domains under this kind of learning typically have the same category of classes as source domains [11]. Due to this, it benefits from exploiting the labels of source domains to learn about multiple related target domains[12, 43]. Some work has a slight domain shift between source and target data like different camera views [44]. At the same time, recent papers have worked on large domain shifts like converting targets into sketch or art domains [12, 45]. Transfer learning is related to MDL or domain adaptation, but focuses on how to generalize better on target tasks [13, 14, 46]. Most of the work in this field uses the popular Imagenet as a source dataset to learn feature representation and learn to transfer to target datasets. The method proposed in [47] uses a pretrained (multi-task) teacher network and decomposes it into multiple task/knowledge-specific factor networks that are disentangled from one another. This factorization leads to sub-networks that can be fine-tuned to downstream tasks, but they rely on knowledge transfer from a teacher network that is pretrained for multiple tasks. Modular deep learning methods [48] focus on transfer learning by avoiding negative task-interference and having parameter-efficient modules.
@@
Factorization methods in MDL/MTL. The method in [49] proposed a unified framework for MTL/MDL using semantic descriptors, but the focus is not on parameter-efficient adaptation. [50] performs MTL/MDL by factorizing each layer in the network after incorporating task-specific information along a separate dimension. Both the networks in [49] and [50] need to be retrained from scratch for new tasks/domains. In contrast, FTN can incrementally learn low-rank factors to add new tasks/domains. [51] proposed a new parameter-efficient network as a replacement for residual networks by incorporating factorized tensors. The results in [51] are limited to learning single-task networks, where the network is only compressed by up to $\sim 60\%$. In [52], the authors proposed a network for MDL using Tucker decomposition.
@@
Transformer-based methods in MDL/MTL There is some work in transformer-based methods. LoRA [53] is a low-rank adaptation method proposed for large language models, which freezes the pre-trained weights of the model and learns low-rank updates for each transformer layer. It updates weight matrices for query and value in every attention layer. Similarly, KAdaptation [54] proposes a parameter-efficient adaptation method for vision transformers. It represents the updates of MHSA layers using the summation of Kronecker products between shared parameters and low-rank task-specific parameters. We compared both of these methods and have shown that FTN outperforms along the number of parameters. Scaling and shifting your features (SSF) [55] is another transformer method for parameter-efficient adaptation that applies element-wise multiplication and addition to tokens after different operations. SSF in principle is similar to fine-tuning the BathNormzation layer in convolutional layers, which has scaling and shifting trainable parameters. FTN trains the Batch Normalization layers and therefore has the same effect as scaling and shifting features when adapting to new tasks. [56] proposed inverted-pyramid multi-task transformer, performs cross-task interaction among spatial features of different tasks in a global context. The DeMT [57] proposes a deformable mixer encoder and task-aware transformer decoder. The proposed encoder leverages channel mixing and deformable convolution operation for informative features while the transformer decoder captures the task interaction.
In summary, our proposed method (FTN) offers a parameter-efficient approach to achieve performance comparable to or better than existing adaptation methods by utilizing a fraction of additional parameters. We demonstrated that, unlike other methods, easy domains do not require the transformation of the backbone by the same complex module, and we can choose the flexibility of the task/domain-specific parameter for a given task/domain.
@@ 3.1 FTN applied to Convolutional layers
In our proposed method, we use task/domain-specific low-rank tensors to adapt every convolutional layer of a pretrained backbone network to new tasks and domains. Let us assume the backbone network has $L$ convolution layers that are shared across all task/domains. We represent the shared network weights as $\mathcal{W}_{\text{shared}}=\{\mathbf{W}_{1},\ldots,\mathbf{W}_{L}\}$ and the low-rank network updates for task/domain $t$ as $\Delta\mathcal{W}_{t}=\{\Delta\mathbf{W}_{1,t},\ldots,\Delta\mathbf{W}_{L,t}\}$. To compute features for task/domain $t$, we update weights at every layer as $\mathcal{W}_{\text{shared}}+\Delta\mathcal{W}_{t}=\{\mathbf{W}_{1}+\Delta\mathbf{W}_{1,t},\ldots,\mathbf{W}_{L}+\Delta\mathbf{W}_{L,t}\}$.
To keep our notations simple, let us only consider $l$th convolution layer that has $k\times k$ filters, $C_{in}$ channels for input feature tensor, and $C_{out}$ channels for output feature tensor. We represent the corresponding $\mathbf{W}_{l}$ as a tensor of size $k^{2}\times C_{in}\times C_{out}$. We represent the low-rank tensor update as a summation of $R$ rank-1 tensors as
@@
Apart from low-rank tensor update, we also optimize over batchnorm layers (BN) for each task/domain [58, 59]. The BN layer learns two vectors $\Gamma$ and $\beta$, each of length $C_{out}$. The BN operation along $C_{out}$ dimension can be defined as element-wise multiplication and addition:
@@
We represent the output of $l$th convolution layer for task/domain $t$ as
@@
Initialization. In our approach, the initialization of the low-rank parameter layers and the pre-trained weights of the backbone network plays a crucial role due to their sensitivity towards performance. To establish a favorable starting point, we adopt a strategy that minimizes substantial modifications to the frozen backbone network weights during the initialization of the task-specific parameter layers. To achieve this, we initialize each parameter layer from the xavier uniform distribution [60], thereby generating $\Delta\mathbf{W}$ values close to 0 before their addition to the frozen weights. This approach ensures the maintenance of a similar initialization state to the frozen weights at iteration 0.
To acquire an effective initialization for our backbone network, we leverage the pre-trained weights obtained from ImageNet. We aim to establish a robust and capable feature extractor for our specific task by incorporating these pre-trained weights into our backbone network.
@@
Effect of batch normalization. In our experiment section, under the ‘FC and BN only’ setup, we have shown that having task-specific batchnorm layers in the backbone network significantly affects the performance of a downstream task/domain. For all the experiments with our proposed approach, we include batch normalization layers as task-specific along with low-rank tensors and classification/decoder layer.

### v2

@@ Abstract
Multi-task and multi-domain learning methods seek to learn multiple tasks/domains, jointly or one after another, using a single unified network. The primary challenge and opportunity lie in leveraging shared information across these tasks and domains to enhance the efficiency of the unified network. The efficiency can be in terms of accuracy, storage cost, computation, or sample complexity. In this paper, we introduce a factorized tensor network (FTN) designed to achieve accuracy comparable to that of independent single-task or single-domain networks, while introducing a minimal number of additional parameters. The FTN approach entails incorporating task- or domain-specific low-rank tensor factors into a shared frozen network derived from a source model. This strategy allows for adaptation to numerous target domains and tasks without encountering catastrophic forgetting. Furthermore, FTN requires a significantly smaller number of task-specific parameters compared to existing methods. We performed experiments on widely used multi-domain and multi-task datasets. We show the experiments on convolutional-based architecture with different backbones and on transformer-based architecture. Our findings indicate that FTN attains similar accuracy as single-task or single-domain methods while using only a fraction of additional parameters per task. The code is available at https://doi.org/10.24433/CO.7519211.v2.
{IEEEkeywords}
Low-rank Adaptation, Multi-domain/Multi-task learning, Tensor Decomposition
@@ 1 Introduction
The primary objective in multi-task learning (MTL) is to train a single model that learns multiple related tasks, either jointly or sequentially. Multi-domain learning (MDL) aims to achieve the same learning objective across multiple domains. MTL and MDL techniques seek to improve overall performance by leveraging shared information across multiple tasks and domains. On the other hand, single-task or single-domain learning does not have that opportunity. Likewise, the storage and computational cost associated with single-task/domain models quickly grows as the number of tasks/domains increases. In contrast, MTL and MDL methods can use the same network resources for multiple tasks/domains, which keeps the overall computational and storage cost small [45, 37].
@@
where $\mathbf{F}_{t}$ represents a function for task/domain $t$ that maps input $\mathbf{x}$ to output $\mathbf{y}_{t}$. We further assume that $\mathbf{F}$ represents a network with a fixed architecture and $\mathcal{W}_{t}$ and $h_{t}$ represent the parameters for task/domain-specific feature extraction and classification/inference heads, respectively. The function in (1) can represent the network for specific task/domain $t$ using the respective $\mathcal{W}_{t},h_{t}$. In the case of MTL, with $T$ tasks, we can have $T$ outputs $\mathbf{y}_{1},\ldots,\mathbf{y}_{T}$ for a given input $\mathbf{x}$. In the case of MDL, we usually have a single output for a given input, conditioned on the domain $t$. Our goal is to learn the $\mathcal{W}_{t},h_{t}$ for all $t$ that maximize the performance of MTL/MDL with minimal computation and memory overhead compared to single-task/domain learning.
@@
Figure 1(a),(b),(c) illustrate three typical approaches for MTL/MDL. First, we can start with a pre-trained network and fine-tune all the parameters $(\mathcal{W}_{t})$ to learn a target task/domain, as shown in Figure 1(a). Fine-Tuning approaches can transfer some knowledge from the pretrained network to the target task/domain, but they effectively use an independent network for every task/domain [24, 45]. Second, we can reduce the parameter and computation complexity by using a completely shared Feature-Extractor (i.e., $\mathcal{W}_{t}=\mathcal{W}_{\text{shared}}$ for all $t$) and learning task/domain-specific heads as last layers, as shown in Figure 1(b). While such approaches reduce the number of parameters, they often result in poor overall performance because of limited network capacity and interference among features for different tasks/domains [24, 45, 52]. Third, we can divide the network into shared and task/domain-specific parameters or pathways, as shown in Figure 1(c). Such an approach can increase the network capacity, provide interference-free paths for task/domain-specific feature extraction, and enable knowledge sharing across the tasks/domains. In recent years, a number of such methods have been proposed for MTL/MDL [45, 16, 29]. While existing methods can provide performance comparable to single-task/domain learning, they require a significantly large number of additional parameters.
In this paper, we propose a parameter-efficient approach to factorize a network into two distinct modules: a shared frozen module and a task/domain-specific module. We refer to this architecture as a factorized tensor network (FTN). FTNs adapt a network to target domains or tasks by learning low-rank tensors and normalization layers, such as batch normalization. An illustration of our proposed method is shown in Figure 1(d), where we represent network parameters as $\mathcal{W}_{t}=\mathcal{W}_{\text{shared}}+\Delta\mathcal{W}_{t}$, where $\Delta\mathcal{W}_{t}$ is a low-rank tensor.
Similar parameter-efficient methods such as [13, 12], use low-rank matrix adaptations to fine-tune their network. Our proposed method represents low-rank adaptations as a summation of R rank-1 tensors, significantly reducing the number of parameters in our network while achieving better performance. LoRA [13] explores a similar approach by using low-rank matrix factorization to adapt networks. However, unlike LoRA, the low-rank tensor factorization in FTN enables greater parameter reduction. Our experiments demonstrate that FTN achieves better results than LoRA. While LoRA was originally designed for transformer architectures, we have shown a natural extension of FTN to convolutional architectures. The recent method SVFT [23] updates weights as a sparse combination of outer products of singular vectors, training only the coefficients of these sparse combinations. Our experiments indicate that FTN achieves superior performance using fewer parameters than SVFT. FTN leverages tensor factorization to efficiently approximate multi-dimensional data. Our main motivation is to exploit the ability of tensor factorization to model complex interactions and dependencies more effectively than traditional 2D matrix representations [35, 18].
A prior work, TAPS [45], differentially learns which layers of a pre-trained network to adapt for a downstream task/domain by learning an indicator function. The network uses adapted weights instead of pre-trained weights if the indicator score is above a certain threshold. This typically involves adapting high-parameterized layers closer to the classifier/head, which uses significantly more parameters than our FTN method. Existing parameter-efficient MTL/MDL methods [24, 25, 21] introduce small task/domain-specific parameters while others [52, 11] add many parameters to boost the performance irrespective of the task complexity. In our work, we demonstrate the flexibility of FTNs by selecting the rank according to the complexity of the task. Other approaches like RCM [16] adapt incrementally to new tasks by reparameterizing the convolutional layer into task-shared and task-specific parameters. However, unlike FTN this architecture shows limitations in adapting based on the complexity of the tasks and performs subpar along performance and parameters. We demonstrate the effectiveness of our method using different MTL and MDL datasets.
@@
The main contributions of this paper are as follows.
  • We propose a new method for MTL and MDL, called factorized tensor networks (FTN), that adds task/domain-specific low-rank tensors to shared weights. FTNs can achieve similar performance as the single-task/domain methods while using a fraction of additional parameters.
  • Our proposed method utilizes tensor-factorization and demonstrates superior parameter-efficiency compared to matrix factorization methods such as LoRA [13] or indicator based adaptation methods such as TAPS [45].
@@
  • We performed empirical analysis to show that the FTNs enable flexibility by allowing us to vary the rank of the task-specific tensors based on the problem complexity.
@@ 2 Related Work
Multi-task learning (MTL) methods commonly leverage shared and task-specific layers in a unified network to solve related tasks [55, 54]. These methods learn shared and task-specific representation through their respective modules. Optimization-based methods [6, 7] devise a principled way to evaluate gradients and losses in multi-task settings. Branched and tree-structured MTL methods [53] enable different tasks to share branches along a tree structure for several layers. Multiple tasks can share computations and features in any layer only if they belong to the same branch in all the preceding layers. [16, 27] proposed MTL networks that incrementally learn new tasks. ASTMT [27] proposed a network that emphasizes or suppresses features depending on the task at hand. RCM [16] reparameterizes the convolutional layer into non-trainable and task-specific trainable modules. We compare our proposed method with these incrementally learned networks. Adashare [40] is another related work in MTL that jointly learns multiple tasks. It learns task-specific policies and network pathways [15].
Multi-domain learning (MDL) focuses on adapting one network to multiple unseen domains or tasks. MDL setup trains models on task-specific modules built upon the frozen backbone network. This setup helps MDL networks avoid negative transfer learning or catastrophic forgetting, which is common among multi-task learning methods. The work by [36, 37] introduces the task-specific parameters called residual adapters. The architecture introduces these adapters as a series or parallel connection on the backbone for a downstream task. Inspired by pruning techniques, Packnet [25] learns on multiple domains sequentially on a single task to decrease the overhead storage, which comes at the cost of performance. Similarly, the Piggyback [24] method uses binary masks as the module for task-specific parameters. These masks are applied to the weights of the backbone to adapt them to new domains. To extend this work, WTPB [26] uses the affine transformations of the binary mask on their backbone to extend the flexibility for better learning. BA2 [1] proposed a budget-constrained MDL network that selects the feature channels in the convolutional layer. It gives a parameter-efficient network by dropping the feature channels based on budget but at the cost of performance. DA3 [46] introduces a memory- and parameter-efficient method with a specific focus on on-device applications. DA3 freezes multiplicative weights and masks and only updates the additive bias terms. [56] paper learns the adapter modules and the plug-in architecture of the modules using NAS. Spot-Tune [11] learns a policy network, which decides whether to pass each image through Fine-Tuning or pretrained networks. It neglects the parameter efficiency factor and emphasises more on performance. TAPS [45] adaptively learns to change a small number of layers in a pretrained network for the downstream task.
@@
Domain adaptation and transfer learning. The work in this field usually focuses on learning a network from a given source domain to a closely related target domain. The target domains under this kind of learning typically have the same category of classes as source domains [41]. Due to this, it benefits from exploiting the labels of source domains to learn about multiple related target domains[43]. Some work has a slight domain shift between source and target data, like different camera views [38]. At the same time, recent papers have worked on significant domain shifts like converting targets into sketch or art domains [43, 57]. Transfer learning is related to MDL or domain adaptation but focuses on better generalizing target tasks [30]. Most of the work in this field uses the popular ImageNet as a source dataset to learn feature representation and learn to transfer to target datasets. The method proposed in [47] uses a pretrained (multi-task) teacher network and decomposes it into multiple task/knowledge-specific factor networks that are disentangled from one another. This factorization leads to sub-networks that can be fine-tuned to downstream tasks, but they rely on knowledge transfer from a teacher network that is pretrained for multiple tasks. Modular deep learning methods [33] focus on transfer learning by avoiding negative task interference and having parameter-efficient modules.
@@
Factorization methods in MDL/MTL. The method in [48] proposed a unified framework for MTL/MDL using semantic descriptors, without focusing on parameter-efficient adaptation. [49] performs MTL/MDL by factorizing each layer in the network after incorporating task-specific information along a separate dimension. Both the networks in [48] and [49] require retraining from scratch for new tasks/domains. In contrast, FTN can incrementally learn low-rank factors to add new tasks/domains. [5] proposed a new parameter-efficient network to replace residual networks by incorporating factorized tensors. The results in [5] are limited to learning single-task networks, where the network is only compressed by up to $\sim 60\%$. In [3], the authors proposed a network for MDL using Tucker decomposition. [51] paper focuses on solving inverse problems in computational imaging applications. The method proposes to modulate the weights of an unrolled pre-trained network for adaptation to multiple domains, measurement models, and noise. The multiplicative modulation is applied on DCNN (a small parameter network) with only rank-1 tensors.
@@
Transformer-based methods in MDL/MTL. COMPACTER [17] is a parameter-efficient fine-tuning method designed for large-scale language models. It inserts task-specific weight matrices into a pretrained model’s weights as a sum of Kronecker products between shared low-rank ”slow” weights and task-specific ”fast” rank-one matrices. Adaptformer [4] introduces an effective adapter-based approach for parameter-efficient fine-tuning of vision transformers for a large variety of downstream visual recognition tasks. The core idea is to insert the lightweight bottleneck adapters into the feed-forward layer of a pretrained transformer. The adapter involves two fully connected layers, a non-linear activation function, and a scaling factor. LoRA [13] is a low-rank adaptation method proposed for large language models, which freezes the pre-trained weights of the model and learns low-rank updates for each transformer layer. It updates weight matrices for query and value in every attention layer. Similarly, KAdaptation [12] proposes a parameter-efficient adaptation method for vision transformers. It represents the updates of MHSA layers using the summation of Kronecker products between shared parameters and low-rank task-specific parameters. We compared both of these methods and have shown that FTN outperforms along the number of parameters. Scaling and shifting your features (SSF) [22] is another transformer method for parameter-efficient adaptation that applies element-wise multiplication and addition to tokens after different operations. SSF, in principle, is similar to fine-tuning the Batch Normalization layer in convolutional layers, which has scaling and shifting trainable parameters. FTN trains the Batch Normalization layers and has the same effect as scaling and shifting features when adapting to new tasks. [50] proposed inverted-pyramid multi-task transformer, performs cross-task interaction among spatial features of different tasks in a global context. Our method, FTN, shares some high-level similarities with other parameter-efficient adaptation methods such as LoRA, as both approaches aim to introduce low-rank factors to adapt networks for multiple tasks and domains. Our method is a natural extension to higher-order tensors, and we demonstrate its effectiveness across both transformer and convolutional network architectures. In addition, our method adds parameter and performance efficiency compared to related methods, as shown by our experiments.
In summary, our proposed method (FTN) offers a parameter-efficient approach to achieve performance comparable to or better than existing adaptation methods by utilizing a fraction of additional parameters. Our primary design consideration was to achieve efficient adaptation, enabling incremental learning with additive factors. To achieve parameter efficiency, we introduce a small number of trainable parameters through low-rank factorization applicable to both convolutional and transformer-based networks. We utilize frozen and trainable task-specific parameters to support incremental learning without forgetting prior knowledge.
@@ 3 Technical Details
Notations. In this paper, we denote scalars, vectors, matrices and tensors by $w$, $\mathbf{w}$, $\mathrm{W}$, and $\mathbf{W}$, respectively. The collective set of tensors (network weights) is denoted as $\mathcal{W}$.
@@ 3.1 FTN applied to Convolutional layers
In our proposed method, we use task/domain-specific low-rank tensors to adapt every convolutional layer of a pretrained backbone network to new tasks and domains. Let us assume the backbone network has $L$ convolutional layers that are shared across all task/domains. We represent the shared network weights as $\mathcal{W}_{\text{shared}}=\{\mathbf{W}_{1},\ldots,\mathbf{W}_{L}\}$ and the low-rank network updates for task/domain $t$ as $\Delta\mathcal{W}_{t}=\{\Delta\mathbf{W}_{1,t},\ldots,\Delta\mathbf{W}_{L,t}\}$. To compute features for task/domain $t$, we update weights at every layer as $\mathcal{W}_{\text{shared}}+\Delta\mathcal{W}_{t}=\{\mathbf{W}_{1}+\Delta\mathbf{W}_{1,t},\ldots,\mathbf{W}_{L}+\Delta\mathbf{W}_{L,t}\}$.
To keep our notations simple, let us only consider $l$th convolutional layer that has $k\times k$ filters, $C_{in}$ channels for input feature tensor, and $C_{out}$ channels for output feature tensor. We represent the corresponding $\mathbf{W}_{l}$ as a tensor of size $k^{2}\times C_{in}\times C_{out}$. We represent the low-rank tensor update as a summation of $R$ rank-1 tensors as
@@
Apart from low-rank tensor update, we also optimize over Batch Normalization layers (BN) for each task/domain [14, 34]. The BN layer learns two vectors $\Gamma$ and $\beta$, each of length $C_{out}$. The BN operation along $C_{out}$ dimension can be defined as element-wise multiplication and addition:
@@
We represent the output of $l$th convolutional layer for task/domain $t$ as
@@
Initialization. To establish a favorable starting point, we adopt a strategy that minimizes substantial modifications to the frozen backbone network weights during the initialization of the task-specific parameter layers. To achieve this, we initialize each parameter layer from the Xavier uniform distribution [10], thereby generating $\Delta\mathbf{W}$ values close to 0 before their addition to the frozen weights. This approach ensures the initial point of our proposed network closely matches the pretrained network closely.
To acquire an effective initialization for our backbone network, we leverage the pretrained weights obtained from ImageNet. We aim to establish a robust and capable feature extractor for our specific task by incorporating these pretrained weights into our backbone network.
@@
Effect of Batch Normalization. In our experiment section, under the ‘FC and BN only’ setup, we have shown that having task-specific Batch Normalization layers in the backbone network significantly affects the performance of a downstream task/domain. For all the experiments with our proposed approach, we include Batch Normalization layers as task-specific along with low-rank tensors and classification/decoder layer.